Find all real zeros of the function. f(x)=x^3+6x^2-4x-24
Answer by Edwin McCravy(20056)

f(x) = x³+6x²-4x-24
Factor x² out of the first two terms
Factor -4 out of the last two terms

f(x) = x²(x+6)-4(x+6)

Factor (x+6) out of both terms:

f(x) = (x+6)(x²-4)

Factor the second parentheses as the difference of squares:

f(x) = (x+6)(x-2)(x+2)

To find the zeros set the expression = 0

       (x+6)(x-2)(x+2) = 0

Use the zero factor property:

        x+6=0; x-2=0;  x+2=0
         x=-6;   x=2;    x=-2

Edwin